In Germany they came first for the Communists, and

Please solve problem 971.

(a) 53 degree

(b) 11.4 degree
Answer #1

A) for warter n=1.33, so:

\theta _{i}=\theta _{p} = tan^{-1}(1.33)

=53.06^{\circ}

B) From water to glass

\theta _{i}=\theta _{p}=tan^{-1}(\frac{n_{t}}{n_{i}})

\theta _{i}=\theta _{p}=tan^{-1}(\frac{1.5}{1.33})=48.44^{\circ}

The angle between glass and water surface

angle=\theta _{i}{}'-\theta _{t}

\theta _{t}=90^{\circ}-53.06^{\circ}=36.94^{\circ}

angle=48.44^{\circ}-36.94^{\circ} =11.5^{\circ}